# New axisymmetric equilibria with flow from

### v1

@@ Abstract
We construct analytic solutions to the generalized Grad-Shafranov (GS) equation, adopting the general linearizing ansatz for the free-function terms it contains, by expanding the genaralized Solovév solution [Ch. Simintzis, G. N. Throumoulopoulos, G. Pantis and H. Tasso, Phys. Plasmas 8, 2641 (2001)]. On the basis of these solutions, we examine how the genaralized Solovév configuration is modified as the values of the free parameters associated with the additional pressure, poloidal-current and electric-field terms are changed. Thus, a variety of equilibria of tokamak, spherical tokamak and spheromak pertinence are constructed including D-shaped configurations with positive and negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
@@ 1 Introduction
The magnetohydrodynamics axisymmetric equilibrium states are governed by the well-known GS equation, a quasi-linear elliptic partial differential equation for the poloidal magnetic flux-function, containing a couple of arbitrary functions of that flux (surface functions), namely the pressure and the poloidal-current functions. This equation is generally solved numerically under appropriate boundary conditions, and to this end, several codes have been developed, e.g. the HELENA code [1]. Also, to easier gain physical intuition and for benchmarking equilibrium codes, analytic solutions have been constructed to linearized forms of the GS equation. In connection with the present study we first mention the widely employed in plasma confinement studies Solovév solution [2]. This solution, corresponding to constant surface-function terms in the GS equation, describes an up-down symmetric equilibrium with D-shaped magnetic surfaces surrounded by a spontaneously formed separatrix with a couple of X-points (cf. Fig. 1). Also, known is the Maschke-Hernegger solution [3, 4], expressed in terms of the Whittaker functions, corresponding to linear choices of the surface-function terms and describing more realistic equilibria with current-density profiles vanishing on the plasma boundary. Several additional analytic solutions to the GS equation are available, e.g. extending the Solovév solution by polynomial contributions to the homogeneous counterpart of the GS equation in order to construct diverted tokamak equilibria [5]; for constant pressure and linear current-density choices of the free terms in the GS equation [6]; for linear choices of both the pressure and current-density terms by imposing D-shaped boundaries or boundaries with a lower X-point [7, 8, 9]; by differentiation of known separated solutions with respect to the separation parameter in order to get simpler solutions [10]; and in elliptic prolate geometry [11]. Also, for the case of the generic linearized form of the GS equation an analytic solution involving infinite series was derived in [12] and was employed to construct equilibrium configurations pertinent to the tokamak ASDEX-Upgrade.
@@
Extended equilibrium investigations have been conducted to include plasma flows and associated electric fields, which play an important role in the creation of advanced confinement regimes in tokamaks, e.g. the transition of the low to high confinement mode. In particular, for incompressible flows, the generalized GS equation (1) shown below was derived [13, 14], containing five free surface quantities. Equation (1) has an additional electric-field term through the electrostatic potential $\Phi$, associated with the component of plasma velocity non-parallel to the magnetic field. For parallel velocity and arbitrary assignment of the free surface-function terms, an extension of the HELENA code was developed in [15], while the generalized Solovév solution to (1) (given by (5) below), which constitutes a basic ingredient of the present study, was obtained in [14]. Owing to the electric field, the respective equilibrium in addition to a separatrix, similar to that of the usual static Solovév equilibrium, possess a third X-point outside the separatrix (cf. Fig. 11 of [14]). Other analytic solutions to the generalized GS equation of tokamak or space plasma pertinence were obtained by adopting a Solovév-like linearizing ansatz and imposing a diverted boundary [16]; alternative linearizing ansatzes [17, 18, 19]; non-linear forms of the generalized GS equation [20]; employing Lie-point symmetries for linear and nonlinear choices of the free-function terms [21, 22, 23]; and by the method of similarity reduction [24, 25].
@@
Aim of the present work is to construct analytically solutions to the generalized GS equation in its generic linearized form by employing an alternative method. This consists in pursuing solutions as expanded generalized Solovév ones, i.e. as superposition of the generalized Solovév solution and a function to be determined. In addition to the novelty of this method we are particularly interested in examining how the Solovév configuration is modified when the influence of the expanding function becomes stronger by changing the pertinent new free parameters.
@@
In section 2 the generalized GS equation and the generalized Solovév solution are briefly reviewed and the method of solving the generalized GS equation adopting the generic linearized ansatz is presented. Then, analytic solutions to the generic linearized generalized GS equation are constructed in section 3. In section 4 we examine how the Solovév configuration is modified by varying the additional free parameters associated with the pressure and the poloidal current in conjunction with the parallel component of the velocity and electric field, thus obtaining a variety of new configurations. The conclusions are summarized in section 5.
@@ 2 Generic linear axisymmetric equilibria with non-parallel flow
Here $(R,\phi,z)$ are cylindrical coordinate with $z$ corresponding to the axis of symmetry; the poloidal magnetic flux function $u(R,z)$ labels the magnetic surfaces; $M_{p}(u)$ is the Mach function of the poloidal velocity with respect to the poloidal Alfven velocity; $\Phi(u)$ is the electrostatic potential; $\rho(u)$ is the plasma density; $X(u)$ relates to the toroidal magnetic field $B_{\phi}=I/R$ through
$I=\frac{X}{\left(1-M_{p}\right)}-R^{2}\frac{\sqrt{\rho}M_{p}\Phi^{\prime}}{\left(1-M_{p}\right)^{1/2}}\,.$ (2)
The operator $\Delta^{*}$ in (1) is $\Delta^{*}:=\partial^{2}/\partial R^{2}-(1/R)\partial/\partial R+\partial^{2}/\partial z^{2}$ and $\mu_{0}$ is the permeability of free space. For vanishing flow, i.e. $M_{p}=\Phi^{\prime}=0$, the surface function $P_{s}(u)$ coincides with the pressure and Eq. (1) reduces to the GS equation.
Following Ref. [14] we make the following choice of the free surface-function terms in (1)
@@
$\frac{1}{2}\frac{d}{du}\left[\frac{\mu_{0}X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}\,,$
@@
Normalizing as $x:=R/R_{0}$, $y:=z/R_{0}$, $U:=u/u_{0}$, $u_{0}:=P_{s0}R_{0}^{4}/(2U_{0}(1+\delta^{2}))$, Eq. (1) assumes the form
$U_{xx}-\frac{1}{x}U_{x}+U_{yy}-2\epsilon x^{4}+-2(1+\delta^{2})x^{2}+2\epsilon=0\,,$ (4)
and the generalized Solovév solution is given by
@@
Figure 1: The equilibrium configuration of the GS solution (5). The red-dashed curve represents the separatrix with $x_{s1}=\sqrt{\epsilon}$, $x_{s2}=\left(-\epsilon\lambda-3\delta^{2}+\sqrt{3}\sqrt{-\epsilon^{2}\lambda^{2}-2\delta^{2}\epsilon\lambda+3\delta^{4}+\delta^{2}\lambda+4\lambda^{2}}/(2\lambda)\right)^{1/2}$ and $y_{s}=\left[(1-\epsilon)^{2}(2\delta^{2}+\lambda(\epsilon+1))\right]^{1/2}$.
As already mentioned in the Introduction, the outermost closed surface (red-dashed line) is a separatrix consisting of an elliptic outer part and a straight inner part parallel to the axis of symmetry; thus, a couple of X-points are created. Expressions of the X-point coordinates and of the inner and outer points of the configuration on the mid-plane $y=0$ in terms of the parameters $\delta$, $\epsilon$ and $\lambda$ are given in the caption of Fig. 1. The magnetic axis is located at ($x=1,y=0$). The parameter $\epsilon$ determines the compactness of the configuration, i.e. for $\epsilon>0$ the equilibrium is diamagnetic and describes either a tokamak or a spherical tokamak, while for $\epsilon=0$ the inner part of the separatrix touches the axis of symmetry forming a spheromak. The parameter $\delta$ relates to the elongation of the magnetic surfaces in the vicinity of the magnetic axis, i.e. the higher the value of $\delta$, the larger the elongation parallel to the axis of symmetry. For $\lambda=0$ the velocity becomes parallel to the magnetic field, while for $\lambda<0$ a third X-point appears outside the separatrix. When the velocity vanishes ($M_{p}=\lambda=0$), the usual static Solovév solution is recovered. Here we will consider equilibria with $\lambda>0$.
@@
where $\bar{\alpha},\bar{\beta},\bar{\gamma}$ are dimensionless parameters. Setting $G:=2\epsilon$, $F:=-2(1+\delta^{2})$, $E:=-2\lambda$, $\alpha:=\bar{\alpha}(u_{0}/U_{0})$, $\beta:=\bar{\beta}(u_{0}/U_{0})$ and $\gamma:=\bar{\gamma}(u_{0}/U_{0})$ we obtain from (1) the following equation:
@@
Note that the solution to (7) for $\alpha=\beta=\gamma=0$ is given by (5). We now expand about the GS solution (5) as
@@
This is an inhomogeneous linear second-order partial differential equation for the unknown function $W(x,y)$. The general solution consists of the sum of the general solution to the counterpart homogeneous equation plus a particular solution of the complete inhomogeneous equation.
@@ 3 Solutions
where the coefficients $c$ and $\bar{a}_{n}$ can be determined by substituting (19) into (15). Here, we will employ the solution $X_{1}(\xi)$.
@@
where $Ai$ and $Bi$ are the Airy functions of the first and second kind. For $\beta\neq 0$ apart from the $k^{2}$-term, Eq. (15) is identical with (25). Therefore a solution of (25) is:
$\bar{W}_{1}(\xi)=\xi\sum_{n=0}^{\infty}d_{n}\xi^{n}\,,$ (27)
with the coefficients $d_{n}$ being determined by the recurrence relations (3) after setting $k=0$ (and replacing $a_{n}$ with $d_{n}$).
Lastly, we will obtain a particular solution of the inhomogeneous equation (3). Introducing once more the variable $\xi=x^{2}$, Eq. (3) is written as:
@@
In view of the converging series solutions (18) and (27) to the respective homogeneous equations (15) and (25), we pursue a similar in form solution to (3):
@@
    When $\beta=0$, then $\epsilon=0$ and therefore the second of (2) becomes $(1/2)d/du(X^{2}/(1-M_{p}^{2}))=0$. In the absence of flow then (2) implies that the plasma is confined by a vacuum toroidal magnetic field ($\beta_{p}=1$ equilibrium); when the non-parallel flow is there the electric-field term in (2) additionally contributes to the poloidal current.
@@
    When $\epsilon\neq 0$ and $\beta\neq 0$, the last of (3) becomes $\beta(3F+2E+6)+24=0$. In this case all the linear terms in the ansatz (2) can remain finite. The latter relation just acts as a constraint on the parameters $\epsilon$, $\delta$ and $\lambda$. To construct the equilibria of Sec. 4 we treated $\epsilon$ and $\lambda$ as free and
@@
Additional free parameters are $\alpha$, $\beta$, $\gamma$, $a_{0}$, $d_{0}$, $b_{0}$, $D_{1},D_{2}$, $R_{0}$, $P_{s0}$, $U_{0}$ and $k$. Also, it is noted that the series expansion (30) was shown numerically to converge and we kept terms up to the order of $\xi^{20}$ in our calculations.
Summarizing, a solution of the general linearized form of the generalized GS equation (2) consists of a separated solution of its homogeneous counterpart in terms of (11), (14) and (18) plus a particular solution of the form (20), expressed in terms of the series (27) and (30).
Figure 2: Left: different equilibrium boundaries are given as the additional linear-in-$U$, pressure term in the first of (2) varies in connection with different values of $\alpha$. The black continuous curve represents the separatrix of the respective genaralized Solovév spheromak configuration. Right: a set of magnetic surfaces for one of the configuration representd by blue, dot-dashed U-curves are shown together with the respective curves of the genaralized Solovév equilibrium.

### v2

@@ Abstract
We construct analytic solutions to the generalized Grad-Shafranov equation, which incorporates both toroidal and poloidal flows. This is achieved by adopting a general linearizing ansatz for the free-function terms of the equation and expanding the generalized Solov’ev solution [Ch. Simintzis, G. N. Throumoulopoulos, G. Pantis and H. Tasso, Phys. Plasmas 8, 2641 (2001)]. On the basis of these solutions, we examine how the genaralized Solov’ev configuration is modified as the values of the free parameters associated with the additional pressure, poloidal-current and electric-field terms are changed. Thus, a variety of equilibria of tokamak, spherical tokamak and spheromak pertinence are constructed, including D-shaped configurations with positive and negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
@@ 1 Introduction
The magnetohydrodynamics axisymmetric equilibrium states are governed by the well-known Grad-Shafranov (GS) equation, a quasi-linear elliptic partial differential equation for the poloidal magnetic flux-function, containing a couple of arbitrary functions of that flux (surface functions), namely the pressure and the poloidal-current functions. This equation is generally solved numerically under appropriate boundary conditions, and to this end, several codes have been developed, e.g. the HELENA code [1]. Also, to easier gain physical intuition and for benchmarking equilibrium codes, analytic solutions have been constructed to linearized forms of the GS equation. In connection with the present study we first mention the widely employed in plasma confinement studies Solov’ev solution [2]. This solution, corresponding to constant surface-function terms in the GS equation, describes an up-down symmetric equilibrium with D-shaped magnetic surfaces surrounded by a spontaneously formed separatrix with a couple of X-points (cf. Fig. 1). Also, known is the Maschke-Hernegger solution [3, 4], expressed in terms of the Whittaker functions, corresponding to linear choices of the surface-function terms and describing more realistic equilibria with current-density profiles vanishing on the plasma boundary. Several additional analytic solutions to the GS equation are available, e.g. extending the Solov’ev solution by polynomial contributions to the homogeneous counterpart of the GS equation in order to construct diverted tokamak equilibria [5]; for constant pressure and linear current-density choices of the free terms in the GS equation [6]; for linear choices of both the pressure and current-density terms by imposing D-shaped boundaries or boundaries with a lower X-point [7, 8, 9]; by differentiation of known separated solutions with respect to the separation parameter in order to get simpler solutions [10]; and in elliptic prolate geometry [11]. Also, for the case of the generic linearized form of the GS equation an analytic solution involving infinite series was derived in [12] and was employed to construct equilibrium configurations pertinent to the tokamak ASDEX-Upgrade.
Extended equilibrium investigations have been conducted to include plasma flows and associated electric fields, which play an important role in the creation of advanced confinement regimes in tokamaks, e.g. the transition of the low to high confinement mode. Specifically, to include axisymmetric toroidal flows, inherently incompressible because of the symmetry, elliptic PDEs were derived in [13] for either isentropic or isothermal magnetic surfaces and analytic solutions were constructed therein. For compressible flows of arbitrary direction the problem reduces to a set of a PDE coupled with a Bernoulli equation involving the pressure [14, 15, 16]. Depending on the value of the poloidal velocity, this PDE can be either elliptic or hyperbolic, that is, there are three critical transition poloidal velocities, thus forming two elliptic and two hyperbolic regions. In the frame of two-fluid model, the electron and ion surfaces depart from the magnetic surfaces and the problem reduces to a set of three PDEs involving respective flux-function labels [17, 18]. The problem was also investigated within the framework of simplified two-fluid models, such as Hall-MHD that neglects the electron inertia [19, 20] and extended MHD [21, 22, 23]. The impact of pressure anisotropy was examined in [24, 25, 26, 27, 28]. In addition, to include fast particle populations the problem was addressed in the frame of hybrid fluid-kinetic modes [29, 30, 31, 32, 33].
For incompressible MHD flows having a poloidal component the generalized GS equation becomes elliptic (Eq. (3)) [34, 35] and decouples from the Bernoulli equation, which can be employed as a formula for the pressure (Eq. 8). Incompressibility, implying that the density becomes uniform on the magnetic surfaces, is an acceptable approximation for fusion laboratory plasmas for two reasons: First, the poloidal velocities of those plasmas lie well within the first elliptic region. Second, since for typical flows in laboratory fusion plasmas, i.e. for Alfvén Mach numbers on the order of 0.01, density variations on the magnetic surfaces are very low.
@@
Equation (3), which contains five free surface quantities, has an additional electric-field term through the electrostatic potential $\Phi$, associated with the component of plasma velocity non-parallel to the magnetic field. For parallel velocity and arbitrary assignment of the free surface-function terms, an extension of the HELENA code was developed in [36], while the generalized Solov’ev solution to (3) (given by (12)), which constitutes a basic ingredient of the present study, was obtained in [35]. Owing to the electric field, the respective equilibrium in addition to a separatrix, similar to that of the usual static Solov’ev equilibrium, possess a third X-point outside the separatrix (cf. Fig. 11 of [35]). Other analytic solutions to the generalized GS equation of tokamak or space plasma pertinence were obtained by adopting a Solov’ev-like linearizing ansatz and imposing a diverted boundary [37]; alternative linearizing ansatzes [38, 39, 40]; non-linear forms of the generalized GS equation [41]; employing Lie-point symmetries for linear and nonlinear choices of the free-function terms [42, 43, 44]; and by the method of similarity reduction [45, 46].
@@
Aim of the present work is to construct analytically solutions to the generalized GS equation in its generic linearized form by employing an alternative method. This consists in pursuing solutions as expanded generalized Solov’ev ones, i.e. as superposition of the generalized Solov’ev solution and a function to be determined. In addition to the novelty of this method we are particularly interested in examining how the Solov’ev configuration is modified when the influence of the expanding function becomes stronger by changing the pertinent new free parameters.
@@
In section 2 the generalized GS equation and the generalized Solov’ev solution are briefly reviewed and the method of solving the generalized GS equation adopting the generic linearized ansatz is presented. Then, analytic solutions to the generic linearized generalized GS equation are constructed in section 3. In section 4 we examine how the Solov’ev configuration is modified by varying the additional free parameters associated with the pressure, the poloidal current in conjunction with the parallel component of the velocity, and electric field, thus obtaining a variety of new configurations. The conclusions are summarized in section 5.
@@ 2 Generic linear axisymmetric equilibria with non-parallel flow
$(1-M_{p}^{2})\Delta^{\star}\psi-\frac{1}{2}\frac{dM_{p}^{2}}{d\psi}|\nabla\psi|^{2}+\frac{1}{2}\frac{d}{d\psi}\left(\frac{X^{2}}{1-M_{p}^{2}}\right)+\mu_{0}R^{2}\frac{dP_{s}}{d\psi}+\mu_{0}\frac{R^{4}}{2}\frac{d}{d\psi}\left[\frac{\rho}{1-M_{p}^{2}}\left(\frac{d\Phi}{d\psi}\right)^{2}\right]=0\,.$ (1)
Here, $(z,R,\phi)$ are cylindrical coordinates with $z$ corresponding to the axis of symmetry; $\psi(R,z)$ is the poloidal magnetic flux function which labels the magnetic surfaces; $M_{p}(\psi)$ is the Mach function of the poloidal velocity with respect to the poloidal-magnetic-field Alfvén velocity; $\rho(\psi)$ and $\Phi(\psi)$ are the density and the electrostatic potential; $X(\psi)$ relates to the toroidal magnetic field and to the poloidal current (cf. relations (4) and (7)); for vanishing flow the surface function $P_{s}(\psi)$ coincides with the pressure (cf. (8)); $\Delta^{\star}=R^{2}\nabla\cdot(\nabla/R^{2})$ and $\mu_{0}$ is the permeability of free space.
Equation (1) can be simplified by the transformation
[…]
which reduces (1) to
@@
Note that no quadratic term as $|{\bf\nabla}u|^{2}$ appears anymore in (3). The transformation (2) does not affect the shape of magnetic surfaces; it just relabels them in terms of the function $u$. It is noted here that there is a misprint of the respective $R^{4}$-term in Eq. (17) of [35], i.e. the density factor should be linear instead of quadratic. To solve Eq. (3), the surface-quantity terms ($(1/2)(d/du)\left[X^{2}/(1-M_{p}^{2})\right]$, $\mu_{0}R^{2}dP_{s}/du$ and $\mu_{0}(R^{4}/2)(d/du)\left[\rho(d\Phi/du)^{2}\right]$) should be assigned as functions of $u$. Also, the equilibrium can be completely determined by calculations in the $u$-space; specifically, the magnetic field, velocity, current density, pressure and electric field are given by the relations:
${\mathbf{B}}=I\nabla\phi+(1-M_{p}^{2})^{-1/2}\nabla\phi\times\nabla u\,,$ (4)
with
$I=(1-M_{p}^{2})^{-1/2}\left[\frac{X}{(1-M_{p}^{2})^{1/2}}-\mu_{0}R^{2}\sqrt{\frac{\rho}{\mu_{0}}}M_{p}\frac{d\Phi}{du}\right]\,,$ (5)
${\mathbf{v}}=\frac{M_{p}}{\sqrt{\mu_{0}\rho}}{\mathbf{B}}-R^{2}(1-M_{p}^{2})^{-1/2}\frac{d\Phi}{du}\nabla\phi\,,$ (6)
${\mathbf{j}}=\frac{1}{\mu_{0}}\left\{(1-M_{p})^{-1/2}\left[\Delta^{\star}u-\frac{1}{2(1-M_{p}^{2})}\frac{d(1-M_{p}^{2})}{du}|\nabla u|^{2}\right]\nabla\phi-\nabla\phi\times\nabla I\right\}\,,$ (7)
$P=P_{s}-\rho\left[\frac{v^{2}}{2}-R^{2}\left(\frac{d\Phi}{du}\right)^{2}\right]\,,$ (8)
${\mathbf{E}}=-\frac{d\Phi}{du}\nabla u$ (9)
with $v^{2}={\mathbf{v}}\cdot{\mathbf{v}}$.
Following Ref. [35] we first make the following choice of the free surface-function terms in (3)
@@
$\frac{1}{2}\frac{d}{du}\left[\frac{X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}\,,$
@@
It is noted that a couple of the five surface quantities involved in the generalized GS equation remain free. Normalizing as $x:=R/R_{0}$, $y:=z/R_{0}$, $U:=u/u_{0}$ with $u_{0}:=P_{s0}R_{0}^{4}/(2U_{0}(1+\delta^{2}))$, Eq. (3) assumes the form
$U_{xx}-\frac{1}{x}U_{x}+U_{yy}-2\lambda x^{4}-2(1+\delta^{2})x^{2}+2\epsilon=0\,,$ (11)
and the generalized Solov’ev solution is given by
@@
Figure 1: The equilibrium configuration of the GS solution (12) on the poloidal plane $x-y$, with $x$ corresponds to the horizontal axis and $y$ to the vertical axis. The red-dashed curve represents the separatrix with $x_{s1}=\sqrt{\epsilon}$, $x_{s2}=\left(-\epsilon\lambda-3\delta^{2}+\sqrt{3}\sqrt{-\epsilon^{2}\lambda^{2}-2\delta^{2}\epsilon\lambda+3\delta^{4}+\delta^{2}\lambda+4\lambda^{2}}/(2\lambda)\right)^{1/2}$ and $y_{s}=\left[(1-\epsilon)^{2}(2\delta^{2}+\lambda(\epsilon+1))\right]^{1/2}$.
As already mentioned in section 1, the outermost closed surface (red-dashed line) is a separatrix consisting of an elliptic outer part and a straight inner part parallel to the axis of symmetry; thus, a couple of X-points are created. Expressions of the X-point coordinates and of the inner and outer points of the configuration on the mid-plane $y=0$ in terms of the parameters $\delta$, $\epsilon$ and $\lambda$ are given in the caption of Fig. 1. The magnetic axis is located at ($x=1,y=0$). The parameter $\epsilon$ determines the compactness of the configuration, i.e. for $\epsilon>0$ the equilibrium is diamagnetic and describes either a tokamak or a spherical tokamak, while for $\epsilon=0$ the inner part of the separatrix touches the axis of symmetry forming a spheromak. The parameter $\delta$ relates to the elongation of the magnetic surfaces in the vicinity of the magnetic axis, i.e. the higher the value of $\delta$, the larger the elongation parallel to the axis of symmetry. For $\lambda=0$ the velocity becomes parallel to the magnetic field, while for $\lambda<0$ a third X-point appears outside the separatrix. When the velocity vanishes ($M_{p}=\lambda=0$), the usual static Solov’ev solution is recovered. Here we will consider equilibria with $\lambda>0$.
@@
where $\bar{\alpha},\bar{\beta},\bar{\gamma}$ are dimensionless parameters. The linear terms in $u$ in (2) result in more realistic equilibrium characteristics, e.g. the constant $\epsilon\left[P_{s0}R_{0}^{2}/(U_{0}(1+\delta^{2}))\right]$ in (2) results in monotonically increasing toroidal-current-density profiles while the additional linear in $u$ term, $-\bar{\beta}\epsilon\left[P_{s0}R_{0}^{2}/(U_{0}^{2}(1+\delta^{2})\right]u$, in (2) results in peaked toroidal-current-density profiles. Also, the linear term $\bar{\gamma}\left[\lambda P_{s0}/(U_{0}^{2}(1+\delta^{2}R_{0}^{2}))\right]u$ in (2) produces larger velocity shear. Setting $G:=2\epsilon$, $F:=-2(1+\delta^{2})$, $E:=-2\lambda$, $\alpha:=\bar{\alpha}(u_{0}/U_{0})$, $\beta:=\bar{\beta}(u_{0}/U_{0})$ and $\gamma:=\bar{\gamma}(u_{0}/U_{0})$ as well as substituting (2) into (3) we straightforwardly obtain the following equation:
@@
Note that the solution to (14) for $\alpha=\beta=\gamma=0$ is given by (12). We now will pursue a solution to (14) in the form of an expansion about the GS solution (12), i.e.
@@
This is an inhomogeneous linear second-order PDE for the unknown function $W(x,y)$. The general solution consists of the sum of the general solution to the counterpart homogeneous equation plus a particular solution of the complete inhomogeneous equation.
@@ 3 Solutions
where the coefficients $c$ and $\bar{a}_{n}$ can be determined by substituting (26) into (22). Here, we will employ the solution $X_{1}(\xi)$. By means of (18), (21), (23) and introducing a new parameter $s$ by $k=sl$ with $l=0,1,2,\ldots$, the solution of the homogeneous PDE (17) is written as:
$W(x,y)=\sum^{\infty}_{l}X_{1}(\xi)\left[C_{l}\cos(sly)+D_{l}\sin(sly)\right]$
$=\sum_{l=0}^{\infty}\sum_{n=0}^{\infty}a_{n,sl}\xi^{n}\left[C_{l}\cos(sly)+D_{l}\sin(sly)\right]\,.$ (27)
The values of the free coefficients $C_{l}$ and $D_{l}$ may be properly fixed in order to impose a boundary. However, here since in the Solov’ev equilibrium the boundary (separatrix) is spontaneously formed we will not impose any boundary condition; accordingly, in (3) we will keep only the first two $l$-terms (for $l=0$ and $1$).
@@
where $Ai$ and $Bi$ are the Airy functions of the first and second kind. For $\beta\neq 0$ apart from the $k^{2}$-term, Eq. (22) is identical to (33). Therefore the solution of (33) is already included in (3) for $l=0$.
Finally, we will obtain a particular solution of the inhomogeneous equation (3). Introducing once more the variable $\xi=x^{2}$, Eq. (3) is written as:
@@
In view of the converging series solutions (25) to the homogeneous equations (22) and (33), we pursue a similarly structured solution to (3):
@@
    When $\beta=0$, then $\epsilon=0$ and therefore the second of (2) becomes
    $(1/2)d/du(X^{2}/(1-M_{p}^{2}))=0$. In the absence of flow then (5) implies that the plasma is confined by a vacuum toroidal magnetic field ($\beta_{p}=1$ equilibrium); when the non-parallel flow is there the electric-field term in (5) additionally contributes to the poloidal current.
@@
    When $\epsilon\neq 0$ and $\beta\neq 0$, the last of (3) becomes $\beta(3F+2E+6)+24=0$. In this case all the linear terms in the ansatz (2) can remain finite. This relation just acts as a constraint on the parameters $\beta$, $\delta$ and $\lambda$. In section 4 we intend to examine the impact of the linear in $u$ electric-field term in (2) on the equilibrium by varying the parameter $\beta$; accordingly, leaving $\beta$ free we will employ the constraint to express $\delta$ in terms of $\beta$ and $\lambda$, i.e.
@@
Additional free parameters are $\alpha$, $\beta$, $\gamma$, $a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$, $R_{0}$, $P_{s0}$, $U_{0}$ and $s$. It is recalled here that $R_{0}$ is a reference length determining the $R$-position of the magnetic axis; the parameters $P_{s0}$ and $U_{0}$ are reference parameters associated with the pressure and poloidal magnetic flux; $\delta$ and $\epsilon$ are geometrical (shaping) parameters of the starting generalized Solov’ev equilibrium (12) shown in Fig. 1; $\lambda$ relates to the electric field term in (12) through the ansatz (2); $\alpha$, $\beta$ and $\gamma$ are physical parameters of the current density, pressure and electric field, respectively, in connection to the additional linear terms in the ansatz (2); and $a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$ and $s$ are shaping parameters associated with the new constructed solution. A list of the free parameters is given in Table 1. Also, it is noted that the series expansion (37) was shown numerically to converge and we kept terms up to the order of $\xi^{20}$ in our calculations.
Table 1: List of the free parameters
Parameters Description
$R_{0}$ Reference $R$-position of the magnetic axis
$P_{s0}$, $U_{0}$ Reference pressure and poloidal magnetic flux
$\delta$, $\epsilon$ Shaping of the generalized Solov’ev equilibrium of Fig. (1)
$\lambda$ Parameter related to the electric field term in the
generalized Solov’ev solution (12)
$\alpha$, $\beta$, $\gamma$ Current density, pressure and electric field parameters
in the ansatz (2)
$a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$, $s$ Shaping parameters associated with the new constructed
solution
$\rho_{0}$, $l_{\rho}$, $M_{p0}$, $l_{M_{p}}$ Density and poloidal Mach function parameters (Eq. (40))
In summary, a solution of the general linearized form of the generalized GS equation (2) consists of a separated solution of its homogeneous counterpart in terms of (18), (21) and (25) plus a particular solution of the form (28), expressed in terms of the series (3) for $l=0$ and (37).
Figure 2: Left: different equilibrium boundaries are given on the poloidal plane $x-y$ as the additional linear-in-$U$ pressure term in the first of (2) varies in connection with different values of $\alpha$. The black continuous curve represents the separatrix of the respective genaralized Solov’ev spheromak configuration. Right: a set of magnetic surfaces for one of the configuration representd by blue, dot-dashed U-curves are shown together with the respective curves of the genaralized Solov’ev equilibrium. $U_{b}$ is the value of the bounding $u$-curve (separatrix).